# Statistical Taylor Expansion111

### v1

@@ 4.3 Types of Calculations to Verify
Algorithms of completely different nature with each representative for its category are needed to test the generic applicability of variance arithmetic [1]. An algorithm can be categorized by comparing the amount of its input and output data [1]:
@@
A transformation algorithm has about equal amounts of input and output data. The information contained in the data remains about the same after transforming. For reversible transformations, a unique requirement for variance arithmetic is to recover the original input uncertainties after a round-trip transformation which is a forward followed by a reverse transformation. Discrete Fourier Transformation is a typical reversible transforming algorithm, which contains the same amount of input and output data, and its output data can be transformed back to the input data using essentially the same algorithm. A test of variance arithmetic using FFT algorithms is provided in Section 8.
@@
A reduction algorithm has much less output data than input data such as numerical integration and statistical characterization of a data set. Some information of the data is lost while other information is extracted during reducing. Conventional wisdom is that a reducing algorithm generally benefits from a larger input data set [5]. A test of statistical characterization through sampling is provided in Section 5.
@@ 5 Math Library Functions
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $e^{x\pm\delta x}$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Each color represents a $\delta x$. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $\ln(x\pm\delta x)$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Each color represents a $\delta x$. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $(1\pm\delta x)^{c}$, for different $c$ as shown by the x-axis, and for different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Each color represents a $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $\sin(x\pm\delta x)$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Each color represents a $\delta x$. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The values and uncertainties of $\ln(e^{x})-x$ vs $x$, as VarDbl Error and VarDbl Uncertainty in the legend. The result of the same calculation using conventional floating-point library functions is shown as Lib Error in the legend.
@@ 5.1 Exponential
Figure 5 shows that the calculated uncertainties using Formula (2.22) agree very well with the measured value deviations for $e^{x+\delta x}$. As a result, all error deviations are very close to $1$, as $1.0021\pm 0.0095$, even though both the uncertainties and the value error deviations increase exponentially with $x$ and $\delta x$.
@@ 5.2 Logarithm
$\ln(x)$ has a distributional zero at $x=0$ so that all $log(x\pm\delta x)$ are rejected when $1/5<P(x)$. Figure 5 shows that the calculated uncertainties using Formula (2.25) agree very well with the measured value deviations for $\ln(x+\delta x)$, and the result error deviations are very close to $1$, as $0.985\pm 0.091$.
@@ 5.3 Power
Figure 5 shows that the calculated uncertainties of $(1\pm\delta x)^{c}$ using Formula (2.31) agree very well with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations close to $1$, as $0.975\pm 0.135$.
@@ 5.4 Sine
$\sin(x)$ has distributional zeros at $x=\pm\pi/2$. Figure 5 shows that the error deviation for $\sin(x+\delta x)$ is $1$ except when $x=\pm\pi/2$ and $\delta x<10^{-8}$, which agrees with the expected Delta distribution at $x=\pm\pi/2$. In other places, the error deviations are very close to $1$, as $0.999\pm 0.009$.
@@ 5.5 Numerical Errors for Library Functions
The combined numerical error of the library function $e^{x}$ and $\ln(x)$ is calculated as $\ln(e^{x})-x$. Figure 5 shows that using either variance arithmetic or the conventional floating-point library functions results in the same value errors. The uncertainties of variance arithmetic bound the value errors effectively, resulting in an error deviation about $0.409$ when $|x|\leq 1$. It looks suspicious when $|x|>1$ the value error is always 00 in both calculations.
@@ 6.1 Uncertainty Propagation in Matrix Determinant
Let vector $[p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [37]. Let \$$\$[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $[p_{1},p_{2}\dots p_{n}]_{n}$ [37]. Formula (6.1) [37] defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$. The sub-determinant [37] $M_{i,j}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (6.2) [37]. Formula (6.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [37].
@@ 6.2 Adjugate Matrix
The square matrix whose element is $(-1)^{i+j}|\mathbf{M}_{j,i}|$ is defined as the adjugate matrix [37] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$. Let $\mathbf{I}$ be the identical matrix for $\mathbf{M}$. Formula (6.15) and (6.16) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ [37], which only involve addition, subtraction and multiplication.
@@ 6.3 Floating Point Rounding Errors
Figure 6.2 shows that the histograms of the normalized errors is Delta distributed for the input uncertainty $10^{-17}$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. Such Delta distribution also holds when the matrix size is less than $7$, or when the input noise precision is 00. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs. As a result, in Figure 6.2, the distribution becomes Gaussian with a hint of Delta distribution. Such delta-like distribution persists until the input noise precision reaches $10^{-10}$, which is also the transition of the two trends in Figure 6.2.
@@
Figure 6.2 suggests that an histogram of normalized errors is more sensitive to rounding errors than the corresponding error deviation, due to averaging effect of the latter. Histograms of normalized errors can be used as a finger-print for the associated calculation. In case error deviation is not known, a nearly Normal-distributed result histogram may be used to identify ideal coverage.
@@ 6.5 First Order Approximation
Formula (6.21) shows the first order approximation of $|\widetilde{\mathbf{M}}|$ leads to the first order approximation of $\delta^{2}|\mathbf{M}|$. It essentially states that when the input precision is much less than $1$, the determinant $|\mathbf{M}|$ of an imprecise matrix $\mathbf{M}$ can be calculated in variance arithmetic using Formula (6.1) directly.
@@ 7.1 Moving-Window Linear Regression Algorithm [1]
[…]
figure Coefficients of X in Formula (7.4) and (7.8) at a current and the next position in a time series of the $\alpha+\beta\;Y$ fit. It shows that the $\beta_{j+1}$ in Formula (7.4) can be calculated progressively from $\beta_{j}$, as Formula (7.6) and (7.5) in order. It also shows that the $\delta^{2}\beta_{j+1}$ in Formula (7.8) cannot be calculated easily from $\delta^{2}\beta_{j}$.
@@
According to Figure 7.1, in which $H$ takes an example value of 2, the calculation of $(\alpha_{j},\beta_{j})$ can be obtained from the previous values of $(\alpha_{j-1},\beta_{j-1})$, to reduce the calculation of Formula (7.3) and (7.4) into a progressive moving-window calculation of Formula (7.5) and (7.6), respectively [1]
@@ 7.2 Variance Adjustment
The specified input uncertainty is always $0.2$. Normal noises of $0.2$ deviation are added to the slopes, except Normal noises of $2$ deviation are added to the slope for $j=10\dots 19$, during which the actual noises is $10$-fold of the specified value.
@@ 8 Fast Fourier Transformation
8.1 Unfaithful Frequency Response of Discrete Fourier Transformation
Each testing algorithm needs to come under careful scrutiny. One important issue is whether the digital implementation of the algorithm is faithful to the original analytic algorithm. For example, the discrete Fourier transformation is only faithful for Fourier transformation at certain frequencies, and it has a different degree of faithfulness for other frequencies. This is called the unfaithful frequency response [1] of the discrete Fourier transformation.
For each signal sequence $h[k],k=0,1\dots N-1$, in which $N$ is a positive integer, the discrete Fourier transformation $H[n],n=0,1\dots N-1$ and its reverse transformation is given by Formula (8.1) and (8.2), respectively [13], in which $k$ is the index time and $n$ is the index frequency for the discrete Fourier transformation  [^16] The index frequency and index time are not necessarily related to time unit. The naming is just a convenient way to distinguish the two opposite domains in the Fourier transformation: the waveform domain vs the frequency domain.
@@ 8.1 Unfaithful Frequency Response of Discrete Fourier Transformation
Formula (8.3) is the discrete forward transformation $H[n]$ of a pure sine signal $h[k]=\sin(2\pi kf/N)$ in which $f$ is the signal frequency. The continuous forward transformation of the $h[k]$ is a delta function at $n=\pm f$ with phase $\pi/2$. $H[n]$ is delta-like function at $n=\pm f$ with phase $\pi/2$ only if $f$ is an integer. In other cases, how much the result of discrete Fourier transformation deviates from continuous Fourier transformation depends on how much $f$ deviates from an integer, e.g., when $f$ is exactly between two integers, the phase of the transformation is that of cosine instead of sine according to Formula (8.3). Examples of unfaithful representations of fractional signal frequency by the discrete Fourier transformation are shown in Figure 8.1. The data for Figure 8.1 is generated using SciPy, which are very reproducible using any other math libraries, including MathLab and Mathematica. Figure 8.1 is a practical reproduction of a previous theoretical figure [1].
$\centering H[n]\@add@centering$ $=\sum_{k=0}^{N-1}\sin(2\pi nk/N)\;e^{i2\pi nk/N}=\frac{1}{2i}\left(\sum_{k=0}^{N-1}e^{i2\pi(n+f)\frac{k}{N}}-\sum_{k=0}^{N-1}e^{i2\pi(n-f)\frac{k}{N}}\right)$
=$=\begin{cases}iN/2,&f\text{ is integer}\\ N/\pi,&f\text{ is integer}+1/2\\ \frac{1}{2}\frac{\sin(2\pi f-2\pi\frac{f}{N})+\sin(2\pi\frac{f}{N})-\sin(2\pi f)e^{-i2\pi\frac{n}{N}}}{\cos(2\pi\frac{n}{N})-\cos(2\pi\frac{f}{N})}&\text{otherwise}\end{cases}$ (8.3)
[…]
figureAn unfaithful Discrete Fourier transformation is demonstrated by the spectra of a few sine signals having amplitude of 1 and slightly different frequencies as shown in legends. The x-axis is for index frequency. The y-axis is for intensity, while the y-axis of the embedded picture is for phase.
Due to its width, a frequency component in an unfaithful transformation may interact with other frequency components of the Discrete Fourier spectrum, thus sabotaging the whole idea of using the Fourier Transformation to decompose a signal into independent frequency components. This means that the common method of signal processing in the Fourier space [13][17][35] may generate artifacts due to its uniform treatment of faithful and unfaithful signal components, which probably coexist in reality. Unlike aliasing [13][17][38], unfaithful representation of the discrete Fourier transformation has an equal presence in the whole frequency range so that it cannot be avoided by sampling the original signal differently.
An unfaithful representation arises from the implied assumption of the discrete Fourier transformation. The continuous Fourier transformation has an infinitive signal range so that:
$h(t)\Leftrightarrow H(s):\;\;\;h(t-\tau)\Leftrightarrow H(s)e^{i2\pi s\tau};$ (8.4)
As an analog, the discrete Fourier transformation $G[n]$ of the signal $h[k],k=1\dots N$ can be calculated mathematically from the discrete Fourier transformation $H[n]$ of $h[k],k=0\dots N-1$:
$G[n]=(H[n]+h[N]-h[0])e^{i2\pi n/N};$ (8.5)
Applying Formula (8.4) to Formula (8.5) results in Formula (8.6).
$h[N]=h[0];$ (8.6)
Thus, the discrete Fourier transformation has an implied assumption that the signal $h[k]$ repeats itself periodically outside its region of $[0,N-1]$. For an unfaithful frequency, $h[N-1]$ and $h[N]$ are discontinuous in regard to signal periodicity, resulting in larger peak width, lower peak height, and wrong phase.
The unfaithfulness of the discrete Fourier transformation to the Fourier transformation is a very serious example of modeling errors, but this problem has not been addressed seriously enough previously. Because the discrete Fourier transformation has widest applications in science and engineering, this problem needs some serious attention.
@@ 8.2 FFT (Fast Fourier Transformation)
  • For each output, each input is only used once, so there is no dependency problem when using Formula (2.5) and (2.9) as arithmetic operations.
@@
  • Each step in the forward transformation thus increases the variance by $2$-fold, so that the result uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, the result uncertainty mean decreases with the FFT order $L$ as $\sqrt{1/2}^{L}$. The result uncertainty means for the roundtrip transformations is thus: $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
@@ 8.3 Testing Signals
Only Formula (8.1) and (8.2) with integer $n$ and $k$ will be used.
@@ 8.4 Library Errors
figure The difference between the library $\sin(x)$ and the indexed $\sin(x)$, for all integer input to the indexed sine functions used in the FFT transformations of FFT order $4$. The uncertainties of the $\sin(x)$ values are also displayed, to mark the periodicity of $\pi$.
@@
Using indexed $\sin(x)$ as standard, Figure 8.4 compares the value errors of the library $\sin(x)$, $\cos(x)/\sin(x)$, and $1/tan(x)$ for increasing FFT orders. It shows that the numerical errors of all these library functions grow exponentially with increasing FFT order. $1/tan(x)$ has the largest errors, and it will not be used for FFT transformations. $\cos(x)/\sin(x)$ has larger errors than those of $\sin(x)$.
The difference between the indexed and library sine functions shows the inconsistency in the library sine functions, because the indexed sine functions is just the library sine functions in the range of $[0,\pi/4]$, and should be related to other ranges according to precise mathematical relations. It is difficult to study the numerical errors of library functions using conventional floating-point arithmetic which has no concept of imprecise values. Having identified the nature and the strength of the numerical errors of the library $\sin(x)$ and $\cos(x)/\sin(x)$, variance arithmetic can show their effects using FFT transformations.

### v2

@@ 4.3 Types of Calculations to Verify
Algorithms of completely different nature with each representative for its category are needed to test the generic applicability of variance arithmetic [1]. An algorithm can be categorized by comparing the amount of its input and output data as [1]:
@@
A transformation algorithm has about equal amounts of input and output data. The information contained in the data remains about the same after transforming. For reversible transformations, a unique requirement for statistical Taylor expansion is to recover the original input uncertainties after a round-trip transformation which is a forward followed by a reverse transformation. Discrete Fourier Transformation is a typical reversible transforming algorithm, which contains the same amount of input and output data, and its output data can be transformed back to the input data using essentially the same algorithm. A test of variance arithmetic using FFT algorithms is provided in Section 8.
@@
A reduction algorithm has much less output data than input data such as numerical integration and statistical characterization of a data set. Some information of the data is lost while other information is extracted during reducing. Conventional wisdom is that a reducing algorithm generally benefits from a larger input data set [5]. For example, discrete Fourier transformation was found to be unfaithful for continuous Fourier transformation [1].
@@ 5 Math Library Functions
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $e^{x\pm\delta x}$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $\log(x\pm\delta x)$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $(1\pm\delta x)^{c}$, for different $c$ as shown by the x-axis, and for different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The calculated uncertainties vs the measured value deviations as well as the measured error deviations for $\sin(x\pm\delta x)$, for different $x$ as shown by the x-axis, and different $\delta x$ as shown in the legend. The uncertainties and the value deviations are drawn using the logarithmic y scales on the left side, while the error deviations are drawn using the linear y scales on the right side. Gaussian noises are used to produce the input noise $\delta x$.
@@
figure The values and uncertainties of $\log(e^{x})-x$ vs $x$, as VarDbl Error and VarDbl Uncertainty in the legend. The result of the same calculation using conventional floating-point library functions is shown as Lib Error in the legend.
@@ 5.1 Exponential
Figure 5 shows that the calculated uncertainties using Formula (2.20) agree very well with the measured value deviations for $e^{x+\delta x}$. As a result, all error deviations are very close to $1$, even though both the uncertainties and the value error deviations increase exponentially with $x$ and $\delta x$.
@@ 5.2 Logarithm
$\log(x)$ has a distributional zero at $x=0$ so that all $log(x\pm\delta x)$ are rejected when $1/5<P(x)$. Figure 5 shows that the calculated uncertainties using Formula (2.23) agree very well with the measured value deviations for $\log(x+\delta x)$, and the result error deviations are very close to $1$.
@@ 5.3 Power
Figure 5 shows that the calculated uncertainties of $(1\pm\delta x)^{c}$ using Formula (2.26) agree very well with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations close to $1$.
@@ 5.4 Sine
$\sin(x)$ has distributional poles at $x=\pm\pi/2$. Figure 5 shows that the error deviation for $\sin(x+\delta x)$ is $1$ except when $x=\pm\pi/2$ and $\delta x<10^{-8}$, which agrees with the expected Delta distribution at $x=\pm\pi/2$. In other places, the error deviations are very close to $1$.
@@ 5.5 Numerical Errors for Library Functions
The combined numerical error of the library function $e^{x}$ and $\log(x)$ is calculated as $\log(e^{x})-x$. Figure 5 shows that using either variance arithmetic or the conventional floating-point library functions results in the same value errors. The uncertainties of variance arithmetic bound the value errors effectively, resulting in an error deviation about $0.409$ when $|x|\leq 1$. It looks suspicious when $|x|>1$ the value error is always 00 in both calculations.
@@ 6.1 Uncertainty Propagation in Matrix Determinant
Let vector $[p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [37]. Let \$$\$[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $[p_{1},p_{2}\dots p_{n}]_{n}$ [37]. Formula (6.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [37]. The sub-determinant $M_{i,j}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (6.2) [37]. Formula (6.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [37].
@@ 6.2 Adjugate Matrix
The square matrix whose element is $(-1)^{i+j}|\mathbf{M}_{j,i}|$ is defined as the adjugate matrix[37] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$. Let $\mathbf{I}$ be the identical matrix for $\mathbf{M}$. Formula (6.15) and (6.16) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ [37].
@@ 6.3 Floating Point Rounding Errors
Figure 6.2 shows that the histograms of the normalized errors is Delta distributed for the input uncertainty $10^{-17}$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. Such Delta distribution also holds when the matrix size is less than $7$, or when the input is precise. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs. As a result, in Figure 6.2, the distribution becomes Gaussian with a hint of Delta distribution. Such delta-like distribution persists until the input noise precision reaches $10^{-10}$, which is also the transition of the two trends in Figure 6.2.
@@
Figure 6.2 suggests that the histogram of normalized errors is more sensitive to rounding errors than the corresponding error deviation, due to averaging effect of the latter. Histograms of normalized errors can be used as a finger-print for the associated calculation. In case error deviation is not known, a nearly Normal-distributed histogram may be used to identify ideal coverage.
@@ 6.5 First Order Approximation
Formula (6.21) shows the first order approximation of $|\widetilde{\mathbf{M}}|$ leads to the first order approximation of $\delta^{2}|\mathbf{M}|$. It states that when the input precision is much less than $1$, the determinant $|\mathbf{M}|$ of an imprecise matrix $\mathbf{M}$ can be calculated in variance arithmetic using Formula (6.1) directly.
@@ 7.1 Moving-Window Linear Regression Algorithm [1]
The calculation of $(\alpha_{j},\beta_{j})$ can be obtained from the previous values of $(\alpha_{j-1},\beta_{j-1})$, to reduce the calculation of Formula (7.3) and (7.4) into a progressive moving-window calculation of Formula (7.5) and (7.6), respectively [1]
@@ 7.2 Variance Adjustment
The specified input uncertainty is always $0.2$. Normal noises of $0.2$ deviation are added to the slopes, except Normal noises of $2$ deviation are added to the slope for $j=10\dots 19$, during which the actual noises are $10$-fold of the specified value.
@@ 8 Fast Fourier Transformation
8.1 Discrete Fourier Transformation (DFT)
For each signal sequence $h[k],k=0,1\dots N-1$, in which $N$ is a positive integer, the discrete Fourier transformation $H[n],n=0,1\dots N-1$ and its reverse transformation is given by Formula (8.1) and (8.2), respectively [13], in which $k$ is the index time and $n$ is the index frequency for the discrete Fourier transformation (DFT)  [^14] The index frequency and index time are not necessarily related to time unit. The naming is just a convenient way to distinguish the two opposite domains in the Fourier transformation: the waveform domain vs the frequency domain..
@@ 8.2 FFT (Fast Fourier Transformation)
  • For each output, each input is only used once, so there is no dependency problem when using Formula (2.14), (2.15), (2.16), and (2.17) as arithmetic operations.
@@
  • Each step in the forward transformation thus increases the variance by $2$-fold, so that the result uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, the result uncertainty mean decreases with the FFT order $L$ as $\sqrt{1/2}^{L}$. The result uncertainty mean for the roundtrip transformations is thus $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
@@ 8.3 Testing Signals
To avoid unfaithfulness of discrete Fourier transformation to continuous transformation, only Formula (8.1) and (8.2) with integer $n$ and $k$ will be used.
@@ 8.4 Library Errors
figure The difference between the library $\sin(x)$ and the indexed $\sin(x)$, for all integer input to the indexed sine functions used in the FFT of FFT order $4$. The uncertainties of the $\sin(x)$ values are also displayed, to mark the periodicity of $\pi$.
@@
Figure 8.4 pin-points the deficit of the current $\sin(x)$ and $\tan(x)$ library functions: near $x=n\pi$. It also indicates the deficit of the current forward rounding error study [12]: near $x=n\pi$, and without mathematical relation test.
Using indexed $\sin(x)$ as standard, Figure 8.4 compares the value errors of the library $\sin(x)$, $\cos(x)/\sin(x)$, and $1/tan(x)$ for increasing FFT orders. It shows that the numerical errors of all these library functions grow exponentially with increasing FFT order. $1/tan(x)$ has the largest errors, and it will not be used for FFT. $\cos(x)/\sin(x)$ has larger errors than those of $\sin(x)$.
The difference between the indexed and library sine functions shows the inconsistency in the library sine functions, because the indexed sine functions is just the library sine functions in the range of $[0,\pi/4]$, and should be related to other ranges according to precise mathematical relations. It is difficult to study the numerical errors of library functions using conventional floating-point arithmetic which has no concept of imprecise values. Having identified the nature and the strength of the numerical errors of the library $\sin(x)$ and $\cos(x)/\sin(x)$, variance arithmetic can show their effects using FFT.